# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v1

@@ 4 RoPE Attention Gradient Calculation
In this section, we compute the gradient of RoPE attention.
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.14),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.15),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.18).
@@
and we note that the 1st equality is by Def. 3.13, the 2nd equality is by chain rule, the 3rd equality is due to Part 1, and the 4th equality is because of Def. 3.13.
@@
and we note that the 1st equality is by Def. 3.14, the 2nd equality is by product rule, the 3rd equality is due to $\frac{\mathrm{d}\mathbf{1}_{n}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, the 4th equality is because of Def. 3.13, and 5th equality derives from basic algebra.
@@
and we note that the 1st equality is by Def. 3.15, the 2nd equality is by product rule, the 3rd equality is due to chain rule, the 4th equality is because of previous parts, the 5th and 6th equalities derive from Def. 3.15.
@@
and we note that the 1st equality is by Def. 3.17, the 2nd equality is by $\frac{\mathrm{d}E_{j_{0},i_{0}}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, and the 3rd equality is due to the previous part.
@@
and we note that the 1st equality is by Def. 3.18, the 2nd equality is by chain rule, and the 3rd equality is due to the previous part. ∎
@@ 5.1 Reformulate the Gradient into Its Closed Form
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.14),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.15),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.18).
@@
where the 1st step follows from Lemma 4.1, and all other steps follow from Fact 3.10.
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.14),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.15),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.18), and
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.14),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.15),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.18),
@@
5.2 Time to Get $s(x)$ and $v(y)$ Functions
@@
Pick $s(x)$ and $v(y)$ from Def. 3.15 and Def. 3.16, then it costs $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ time to get $s(x)$, and it costs $\mathcal{T}_{\mathrm{mat}}(n,d,d)$ time to get $v(y)$.
@@ 5.3 Time Complexity for Computing $\ell(x)$ Functions
We have $\ell(x)$ from Def. 3.17, then it costs $\mathcal{T}_{\mathrm{mat}}(n,n,d)+O(nd)$ to calculate $\ell(x)$.
@@ 5.5 Time Complexity for Computing the Gradient of RoPE Attention
Lemma 5.7 (RoPE attention gradient computation time complexity).
@@ 6 Low Rank Approximation of Attention
This section presents the fast running time using the polynomial method and Fast Fourier transform.
@@ 6.1 Approximate $s$ Using Low Rank Approximation
Lemma 6.2.
@@ 6.2 Approximate $\ell$ Using Low Rank Approximation
Lemma 6.3.
Let $d$ equal $O(\log n)$. Suppose we can use $O(\log n)$ bits to write every entry in $E,v(y)\in\mathbb{R}^{n\times d}$. Define the $\ell(x)\in\mathbb{R}^{n\times d}$ as specified in Def. 3.17. Then, we have $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$ such that $\|U_{1}V_{1}^{\top}v(y)-E-\ell(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n).$
@@
where the 1st is because of Def. 3.17, 2nd step is based on the distributive law, and 3rd step is due to Lemma 6.2. ∎
@@ 6.3 Approximate $\beta$ Using Low Rank Approximation
Lemma 6.4.
Let $k_{2}=n^{o(1)}$. We define $\ell(x)\in\mathbb{R}^{n\times d}$ based on Def. 3.17, and $v(y)\in\mathbb{R}^{n\times d}$ based on Def. 3.16. We suppose $\beta(x)$ is equal to $v(y)\ell(x)^{\top}$, which is an $n\times n$ matrix. Let $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ such that $\|U_{2}V_{2}^{\top}-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. In $n^{1+o(1)}$ time, we can get $U_{2},V_{2}$.
@@
Given that all low rank matrices, we have $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ where $k_{2}=\max\{d,k\}+d=n^{o(1)}$.
@@ 6.4 Approximate $\gamma$ Using Low Rank Approximation
Lemma 6.5.
@@
Then, we have the following by Fact 3.9.
@@
where the first inequality is because of the def. of $\gamma_{1}(x)$, the second equality is due to the def. of $U_{3},V_{3}$, the third equality is based on Fact 3.9, the fourth equality is due to the def. of $\widetilde{s}(x)$ and $\widetilde{\beta}(x)$, the fifth equality is due to simple arithmetic, the sixth inequality is because of the triangle inequality, and the seventh inequality derives from Lemma 6.2 and Lemma 6.4. ∎
Lemma 6.6.
@@
Then, we define $\widetilde{\rho}(x)_{j_{0}}$ as the inner product of $\widetilde{s}(x)_{j_{0}}$ and $\widetilde{\beta}(x)_{j_{0}}$, and by Fact 3.10, we have $\widetilde{\rho}(x)_{j_{0}}=(U_{1}V_{1})_{j_{0},*}\cdot(U_{2}V_{2})_{j_{0},*}^{\top}$
@@ 6.5 Fast Computation in Almost Linear Time
Theorem 6.7 (Main result).
@@
We can use low rank approximation technique to represent $\widetilde{\gamma}_{1}(x)=U_{3}V_{3}^{\top}$ and $\widetilde{\gamma}_{2}(x)=U_{3}V_{3}^{\top}$ as the approximation to $\gamma_{1}(x)$ and $\gamma_{2}(x)$ respectively.
@@ 8 Conclusion
This paper presents the first practical and efficient backward gradient algorithm under the assumption of bounded entries for the RoPE-based attention mechanism. By leveraging polynomial methods and the Fast Fourier Transform, we achieve almost linear time complexity, making the forward and backward computations comparably efficient. Additionally, we demonstrate that conditions exist under which performance better than quadratic can be realized, consistent with the lower bounds suggested by the Strong Exponential Time Hypothesis (SETH). These findings not only improve the computational efficiency of RoPE-based attention mechanisms but also provide a foundation for exploring sub-gradient computations in other advanced attention variants of neural networks. This work highlights the connection between algorithm design and computational complexity theory, unveiling new possibilities for the development of large transformer models. Future research could extend these results to cases involving unbounded entries and assess the real-world implications of these theoretical advancements for large language models. Furthermore, applying this approach to other positional encoding mechanisms could further enhance the scalability of state-of-the-art transformer models.

### v2

@@ 4 RoPE Attention Gradient Calculation
In this section, we compute the entry-wise gradient of the RoPE attention loss function from Lemma 3.18
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17).
@@
and we note that the 1st equality is by Def. 3.12, the 2nd equality is by chain rule, the 3rd equality is due to Part 1, and the 4th equality is because of Def. 3.12.
@@
and we note that the 1st equality is by Def. 3.13, the 2nd equality is by product rule, the 3rd equality is due to $\frac{\mathrm{d}\mathbf{1}_{n}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, the 4th equality is because of Def. 3.12, and 5th equality derives from basic algebra.
@@
and we note that the 1st equality is by Def. 3.14, the 2nd equality is by product rule, the 3rd equality is due to chain rule, the 4th equality is because of previous parts, the 5th and 6th equalities derive from Def. 3.14.
@@
and we note that the 1st equality is by Def. 3.16, the 2nd equality is by $\frac{\mathrm{d}E_{j_{0},i_{0}}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, and the 3rd equality is due to the previous part.
@@
and we note that the 1st equality is by Def. 3.17, the 2nd equality is by chain rule and the 3rd equality is due to the previous part. ∎
@@ 5.1 Reformulate the Gradient into Its Closed Form
In this section, we reformulate the entry-wise gradient of the RoPE loss function from Lemma 4.1 into its matrix form.
We first begin with reformulating the gradient with respect to the entire vector $x$.
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17).
@@
where the 1st step follows from Lemma 4.1, and all other steps follow from Fact 3.9.
@@
Next, we show our reformulation of the gradient by dropping the index $i_{0}$ from $\mathsf{Loss}(x)_{j_{0},i_{0}}$
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17), and
@@
Finally, we reformulate the gradient into its matrix form.
@@
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17),
@@
5.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $s(x)$ and $v(y)$.
@@ 5.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
Pick $s(x)$ and $v(y)$ from Def. 3.14 and Def. 3.15, then it costs $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ time to get $s(x)$, and it costs $\mathcal{T}_{\mathrm{mat}}(n,d,d)$ time to get $v(y)$.
@@ 5.3 Time Complexity for Computing $\ell(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $\ell(x)$.
@@
We have $\ell(x)$ from Def. 3.16, then it costs $\mathcal{T}_{\mathrm{mat}}(n,n,d)+O(nd)$ to calculate $\ell(x)$.
@@ 5.4 Time Complexity for Computing $\beta(x)$ and $\gamma(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $\beta(x)$ and $\gamma(x)$.
@@ 5.5 Time Complexity for Computing the Gradient of RoPE Attention
Finally, We are able to present the time complexity for computing the gradient of RoPE attention.
Theorem 5.7 (RoPE attention gradient computation time complexity).
@@ 6 Low Rank Approximation of Attention
This section presents the fast running time using the low-rank approximations where the low-rank matrices are generated from the polynomial method (see Lemma 3.4).
In recent work [7], they present an almost linear-time algorithm to compute forward computations of RoPE attention as follows.
@@ 6.1 Approximate $s$ Using Low Rank Approximation
In this section, we use the low-rank approximation technique to approximate $s(x)$
Lemma 6.2 (Low Rank Approximate $s(x)$).
@@ 6.2 Approximate $\ell$ Using Low Rank Approximation
In this section, we use the low-rank approximation technique to approximate $\ell(x)$
Lemma 6.3 (Low Rank Approximate $\ell(x)$).
Let $d$ equal $O(\log n)$. Suppose we can use $O(\log n)$ bits to write every entry in $E,v(y)\in\mathbb{R}^{n\times d}$. Define the $\ell(x)\in\mathbb{R}^{n\times d}$ as specified in Def. 3.16. Then, we have $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$ such that $\|U_{1}V_{1}^{\top}v(y)-E-\ell(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n).$
@@
where the 1st is because of Def. 3.16, 2nd step is based on the distributive law, and 3rd step is due to Lemma 6.2. ∎
@@ 6.3 Approximate $\beta$ Using Low Rank Approximation
In this section, we use the low-rank approximation technique to approximate $\beta(x)$
Lemma 6.4 (Low Rank Approximate $\beta(x)$).
Let $k_{2}=n^{o(1)}$. We define $\ell(x)\in\mathbb{R}^{n\times d}$ based on Def. 3.16, and $v(y)\in\mathbb{R}^{n\times d}$ based on Def. 3.15. We suppose $\beta(x)$ is equal to $v(y)\ell(x)^{\top}$, which is an $n\times n$ matrix. Let $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ such that $\|U_{2}V_{2}^{\top}-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. In $n^{1+o(1)}$ time, we can get $U_{2},V_{2}$.
@@
Given that all low-rank matrices, we have $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ where $k_{2}=\max\{d,k\}+d=n^{o(1)}$.
@@ 6.4 Approximate $\gamma$ Using Low Rank Approximation
In this section, we use the low-rank approximation technique to approximate $\gamma(x)$. Specifically, we apply the polynomial methods to $\gamma_{1}(x)$ and $\gamma_{2}(x)$ where $\gamma(x)=\gamma_{1}(x)-\gamma_{2}(x)$.
First, we show the low-rank approximation of $\gamma_{1}(x)$.
Lemma 6.5 (Low Rank Approximate $\gamma_{1}(x)$).
@@
Then, we have the following by Fact 3.8.
@@
where the first inequality is because of the def. of $\gamma_{1}(x)$, the second equality is due to the def. of $U_{3},V_{3}$, the third equality is based on Fact 3.8, the fourth equality is due to the def. of $\widetilde{s}(x)$ and $\widetilde{\beta}(x)$, the fifth equality is due to simple arithmetic, the sixth inequality is because of the triangle inequality, and the seventh inequality derives from Lemma 6.2 and Lemma 6.4. ∎
Next, we show the low-rank approximation of $\gamma_{2}(x)$.
Lemma 6.6 (Low Rank Approximate $\gamma_{2}(x)$).
@@
Then, we define $\widetilde{\rho}(x)_{j_{0}}$ as the inner product of $\widetilde{s}(x)_{j_{0}}$ and $\widetilde{\beta}(x)_{j_{0}}$, and by Fact 3.9, we have $\widetilde{\rho}(x)_{j_{0}}=(U_{1}V_{1})_{j_{0},*}\cdot(U_{2}V_{2})_{j_{0},*}^{\top}$
@@ 6.5 Fast Computation in Almost Linear Time
In this section, we present our main result. With the low-rank approximation, we can approximate the RoPE gradient computations in almost linear time.
Theorem 6.7 (Main result, Low Rank Approximate RoPE Attention Gradient).
@@
We can use low-rank approximation technique to represent $\widetilde{\gamma}_{1}(x)=U_{3}V_{3}^{\top}$ and $\widetilde{\gamma}_{2}(x)=U_{3}V_{3}^{\top}$ as the approximation to $\gamma_{1}(x)$ and $\gamma_{2}(x)$ respectively.
@@ 8 Conclusion
This paper presents the first practical and efficient backward gradient algorithm, assuming bounded entries for the RoPE-based attention mechanism. We achieve almost linear time complexity by leveraging polynomial methods and the Fast Fourier Transform, making the forward and backward computations comparably efficient. Additionally, we demonstrate that conditions exist under which performance better than quadratic can be realized, consistent with the lower bounds suggested by the Strong Exponential Time Hypothesis (SETH). These findings not only improve the computational efficiency of RoPE-based attention mechanisms but also provide a foundation for exploring sub-gradient computations in other advanced attention variants of neural networks. This work highlights the connection between algorithm design and computational complexity theory, unveiling new possibilities for the development of large transformer models. Future research could extend these results to cases involving unbounded entries and assess the real-world implications of these theoretical advancements for large language models. Furthermore, applying this approach to other positional encoding mechanisms could further enhance the scalability of state-of-the-art transformer models.